# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

### v1

@@ 6 Nested normals and facial subspaces
Proof of Lemma 6.5.
It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$. If this statement is shown to hold, applying it recursively generates the sequence of facial subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps. Note that the dimension of the facial subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.6, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$. Note that $\operatorname{\Pi}_{U^{\perp}}(V+S)\subseteq V+S$. This becomes apparent from the following equality: $V+S=(V+S)+U=\operatorname{\Pi}_{U^{\perp}}(V+S)+U,$ where the first equality holds because $U\subseteq V$ implies that $V=V+U$.
@@
Together, Lemma 6.3 and Lemma 6.5 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of facial subspaces, using only nested normals. Indeed, the set of facial subspaces forms a tree, where the root is $\mathbb{R}^{n}$ (note that it is always a facial subpace, and that it contains any other facial subspace), and every edge of the tree is associated with a nested normal. The induction statement is the following.
Proposition 6.1 (Induction by nested normal).
Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If we have show that:
@@
    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$. Then it follows that
    $\forall V\in\mathcal{F}(C;S),\,A(W).$
6.1 Minimal facial subspace
Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$. It is the only face with $x$ in its relative interior. In this section we develop an analogous notion for facial subspaces, and characterize this facial subspace by using nested normals.
Definition 6.5 (The generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated facial subspace $H_{C}(S)$ of $C$, generated by $S\subseteq\mathbb{R}^{n}$, to be the facial subspace that is contained in every other facial subspace in $\mathcal{F}(C;S)$.
@@ 6.1 Minimal facial subspace
Lemma 6.7 (Facial subspaces are closed under intersections).
@@
Theorem 6.3 (Existence and uniqueness of the generated facial subspace).
The generated facial subspace of Definition 6.5 exists and is unique.
Proof of Theorem 6.3.
We argue that there exists a facial subspace which contains no other. Arguing for the existence by contradiction, if every facial subspace of $C$ at $S$ properly contains some other facial subspace of $C$ at $S$, starting at any facial subspace, we can find an infinite chain of facial subspaces with dimension decreasing by one at each step. This is not possible because the dimension is finite. Therefore this chain of subspaces must terminate at some subspace which contains no other.
We now argue that there is only one subspace that contains no other. Let $V_{1},V_{2}$ be two distinct minimal facial subspaces of $C$ at $S$. Then by Lemma 6.7, $V_{1}\cap V_{2}$ is a facial subspace of $C$ at $S$ properly contained in both $V_{1}$ and $V_{2}$, which contradicts $V_{1}$ and $V_{2}$ being minimal.
We have shown the existence and uniqueness of a facial subspace that contains no other. From any facial subspace, we can start a proper containment chain as in our existence argument. The subspace at the end of this chain is the subspace which contains no other, which concludes the proof. ∎
As might be expected, the generated facial subspace is closely related to the generated face.
Lemma 6.8 (Relation of the generated face to the generated facial subspace).
@@
By Theorem 6.1, $H_{C}(S)$ being the smallest facial subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own facial subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a facial subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is an affine facial subspace, as implied by Theorem 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other facial subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
We can now reformulate Definition 4.1 into an equivalent definition.
Definition 6.6 (Equivalent definition of the joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ from Definition 4.1 is
@@
The next result characterizes generated facial subspaces in terms of nested normals.
Lemma 6.9 (The generated facial subspace is the only facial subspace with no nested normals).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated facial subspace $H_{C}(S)$ is the only facial subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
Proof of Lemma 6.9.
The generated facial subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the facial subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Lemma 6.3.
For any facial subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$. Note that $H_{C}(S)\subseteq V$ by Definition 6.5. The inclusion cannot be proper, because otherwise, by Lemma 6.5, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$. ∎
@@ 6.2 Nested normals of a difference of sets
Lemma 6.10 (Characterization of the conditional normals of a difference of sets).
@@
To show Lemma 6.10, we require the following lemma.
@@
We start with the inclusion $(\subseteq)$. Any $v\in N_{C-D}(0)$ has the property that $\forall c\in C,d\in D$, $\langle v,(c-d)-0\rangle\leq 0$. Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$. By a similar argument, $v\in N_{D}(S)$, which shows the inclusion.
@@
Proof of Lemma 6.10.
We argue by structured induction over the facial subspaces of CDC-D, as described in Proposition 6.1. Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
@@
The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a facial subspace. Point 2) is given by Lemma 6.11, and point 3) holds trivially.
For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true. Let $v\in N_{C-D}^{V}(0)\setminus\{0\}$, and $U=V\cap v^{\perp}$. In the rest of the proof, we show that $A(U)$ is true. Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Lemma 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
@@
Considering Definition 4.1 in the light of Corollary 6.1, $T(C,D)=H_{C-D}(0)$. So $T\in\mathcal{F}_{C-D}(0)$, which by Lemma 6.10 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$. By the definition of facial subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$. ∎
@@ 7 Proofs
We first show a smaller result, after which each subsection is dedicated to proving one result from Section 4.
@@
where the first equality follows by Lemma 6.1 and the second by Lemma 6.2. Then by Lemma 6.9, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.8, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
From Proposition 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
@@
7.1 Regularity in the joint facial subspace
We now show the main property of the joint facial subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [10, Theorem 2.19].
Lemma 7.1 (Equivalence of collisions and separation).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $S\subseteq C\cap D$ non-empty, the following statement holds.
@@ 7.1 Regularity in the joint facial subspace
    $\forall s\in S,\langle v,s\rangle=\beta$.
@@
    $\forall c\in C$, $\langle y,v\rangle\geq\beta$.
@@
    $\forall d\in D$, $\langle d,v\rangle\leq\beta$. Then $\forall y\in C,s\in S$, $\langle v,y-s\rangle\geq 0,$ and so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, so $v\in N_{C}(S)\cap(-N_{D}(S))$, which shows that $N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.
@@
Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Lemma 6.9, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Lemma 6.10, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
By applying Lemma 7.1 in the space $T_{a}$, the result follows. Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$. Indeed, $T$ is a finite-dimensional subspace, and thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$. ∎
7.2 First characterization
In this section we prove Proposition 4.1 and Theorem 4.1.
Lemma 7.2 (nested constraint subspaces for nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, $S\subseteq C$ be non-empty, and $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then

### v3

@@ 6 A framework for local face lattices
Proof of Theorem 6.4.
It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$. If this statement is shown to hold, applying it recursively generates the sequence of supporting subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps. Note that the dimension of the supporting subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.3, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$.
Note that $\operatorname{\Pi}_{U^{\perp}}(V+S)\subseteq V+S$. This becomes apparent from the following equality: $V+S=(V+S)+U=\operatorname{\Pi}_{U^{\perp}}(V+S)+U,$ where the first equality holds because $U\subseteq V$ implies that $V=V+U$.
@@
Together, Theorem 6.3 and Theorem 6.4 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of supporting subspaces, using only nested normals. Indeed, the set of supporting subspaces forms a tree, where the nodes are the supporting subspaces, the root is $\mathbb{R}^{n}$ (note that it is always a supporting subspace, and that it contains any other supporting subspace), and every edge of the tree is associated with a nested normal. Crucially, Theorem 6.4 shows that starting from the root, and generating the tree by using the nested normals, we attain all supporting subspaces. Induction on this tree then takes the following form.
Corollary 6.2 (Induction by nested normal).
Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If the following statements hold:
@@
    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$, then it follows that
    $\forall V\in\mathcal{F}(C;S),\,A(V).$
Proof of Corollary 6.2.
Consider any $U\in\mathcal{F}(C;S)$, and let the statement $A$ satisfy both stated conditions. Then by Theorem 6.4, we have supporting subspaces $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ and nested normals $v_{1},\ldots,v_{k}$ as in Theorem 6.4. $A(V_{0})=\mathrm{true}$ because $V_{0}=\mathbb{R}^{n}$. Also, $\forall i\in[k]$, $A(V_{i-1})\implies A(V_{i})$ because of the existence of the nested normal $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ with the property that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$. Then $A(U)$ holds by classical induction, because $U=V_{k}$. ∎
6.1 Minimal supporting subspace
Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$. It is the only face with $x$ in its relative interior. In this section, we develop an analogous notion for supporting subspaces, and characterize this supporting subspace by using nested normals.
Definition 6.5 (The generated supporting subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated supporting subspace $H_{C}(S)$ of $C$, generated by $S$, to be the supporting subspace that is contained in every other supporting subspace in $\mathcal{F}(C;S)$.
@@ 6.1 Minimal supporting subspace
Lemma 6.4 (Supporting subspaces are closed under intersections).
@@
Proposition 6.1.
The generated supporting subspace of Definition 6.5 exists and is unique.
Proof of Proposition 6.1.
For existence, starting at any supporting subspace $U\in\mathcal{F}(C;S)$, either $U$ contains no other supporting subspace in $\mathcal{F}(C;S)$, or we have a $U^{\prime}\in\mathcal{F}(C;S)$ such that $U^{\prime}\subsetneq U$. Repeating this argument, we eventually find a supporting subspace that contains no other, because the dimension is reduced at every step, and the ambient dimension is finite.
For uniqueness, let $V_{1},V_{2}\in\mathcal{F}(C;S)$ be such that they properly contain no other supporting subspace in $\mathcal{F}(C;S)$. Then by Lemma 6.4, $V_{1}\cap V_{2}\in\mathcal{F}(C;S)$ is properly contained in both $V_{1}$ and $V_{2}$, which is a contradiction. ∎
As might be expected, the generated supporting subspace is closely related to the generated face.
Lemma 6.5 (Relation between the generated face and the generated supporting subspace).
@@
By Theorem 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Theorem 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
We can now restate Definition 4.1 with specialized terms.
Definition 6.6 (Reformulated definition of the joint supporting subspace).
The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$, as defined in Definition 4.1, is
@@
The next result characterizes generated supporting subspaces in terms of nested normals.
Theorem 6.5 (The generated supporting subspace is the only supporting subspace with no nested normals).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only supporting subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
Proof of Theorem 6.5.
The generated supporting subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the supporting subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Theorem 6.3.
For any supporting subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$. Note that $H_{C}(S)\subseteq V$ by Definition 6.5. The inclusion cannot be proper, because otherwise, by Theorem 6.4, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$. ∎
@@ 6.2 Nested normals of a difference of sets
Theorem 6.6 (Characterization of the nested normals of a difference of sets).
@@
To show Theorem 6.6, we require the following lemma.
@@
We start with the inclusion $(\subseteq)$. Any $v\in N_{C-D}(0)$ has the property that $\forall c\in C,d\in D$, $\langle v,(c-d)-0\rangle\leq 0$. Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$. By a similar argument, $v\in-N_{D}(S)$, which shows the inclusion.
@@
Proof of Theorem 6.6.
We argue by structured induction over the supporting subspaces of CDC-D, as described in Corollary 6.2. Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
@@
The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a supporting subspace. Point 2) is given by Lemma 6.6, and point 3) holds trivially.
For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true. Let $v\in N_{C-D}^{V}(0)\setminus\{0\}$, and $U=V\cap v^{\perp}$. In the rest of the proof, we show that $A(U)$ is true. Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Theorem 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
@@
By Definition 6.6, $T(C,D)=H_{C-D}(0)$. So $T\in\mathcal{F}_{C-D}(0)$, which by Theorem 6.6 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$. By the definition of supporting subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$. ∎
@@ 7 Proofs
We begin with the proof of Proposition 5.1, after which each subsection is dedicated to one result from Section 4.
@@
where the first equality follows by Lemma 6.1 and the second by Lemma 6.2. Then by Theorem 6.5, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.5, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
From Lemma 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
@@
7.1 Regularity in the joint supporting subspace
We now show the main property of the joint supporting subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [mordukhovich_VariationalAnalysisApplications_2018, Theorem 2.19].
Lemma 7.1 (Separation from local normal cones).
For convex sets $C,D\subseteq\mathbb{R}^{n}$ and $S\subseteq C\cap D$ nonempty, the following equivalence holds:
@@ 7.1 Regularity in the joint supporting subspace
    $\langle v,s\rangle=\beta$ for all $s\in S$;
@@
    $\langle v,c\rangle\geq\beta$ for all $c\in C$;
@@
    $\langle v,d\rangle\leq\beta$ for all $d\in D$.
Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.\}\$. ∎
@@
Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Theorem 6.5, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Theorem 6.6, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
By applying Lemma 7.1 in the space $T_{a}$, the result follows. Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$ because $T$ is a finite-dimensional subspace, thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$. ∎
7.2 Analytical characterizations
In this section we prove Lemma 4.1 and Theorem 4.1.
Lemma 7.2 (Nested normals of constrained set).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $S\subseteq C$ a nonempty set. Let $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then